# WorldSense: Evaluating Real-world Omnimodal Understanding for Multimodal LLMs

### v1

@@ Abstract
In this paper, we introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs. In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality. Based on our WorldSense, we extensively evaluate various state-of-the-art models. The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ( best accuracy). We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
@@ 1 Introduction
The ability to comprehend and reason about multimodal inputs—ranging from visual and textual to auditory, tactile, and beyond—is fundamental for both human and artificial agents to navigate and interpret the world. For example, when driving a car, a human driver integrates visual information (e.g., recognizing road signs, traffic lights, and obstacles), auditory cues (e.g., hearing the honking of another car or a siren approaching from behind), and tactile feedback (e.g., the feel of the steering wheel, the vibrations of the road, or the responsiveness of the brakes) to make real-time decisions and ensure safe navigation. This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and the development of embodied agents designed to interact naturally in the world.
@@
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [49, 24, 48, 57, 59, 75, 43, 12] have led to remarkable progress on a series of tasks, for example, classification [36], captioning [1, 10, 35], question-answering [56, 50, 40], OCR [45, 76], segmentation [25, 67, 22], autonomous driving [46, 52, 4] and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have begun incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench [31] and AV-Odyssey Bench [19] mainly emphasize image evaluation, whereas other benchmarks [18, 29, 70] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
@@
In this paper, we introduce WorldSense, the first benchmark designed to comprehensively evaluate MLLMs’ ability to perceive, understand, and reason through the integration of omni-modal information in real-world contexts. WorldSense is defined by three key features: (i) Collaboration of Omni-Modality: At the core of WorldSense is the emphasis on the interplay between audio and visual information, as illustrated in Figure 1. Each question is designed to require both modalities for a correct answer; the absence of either modality leads to an incorrect response. Such design rigorously tests a model’s ability to simultaneously process multiple sensory inputs for accurate understanding; (ii) Diversity of Videos and Tasks: WorldSense encompasses 1,662 audio-visual synchronized videos across 8 domains and 67 fine-grained subcategories. It includes 3,172 multiple-choice questions spanning 26 distinct cognitive tasks, from basic perception to complex reasoning. This diversity enables a systematic and comprehensive evaluation of MLLMs’ multimodal understanding capabilities; (iii) High-Quality Annotations: To ensure reliability, all QA pairs in WorldSense are manually labeled by 80 expert annotators and undergo multiple rounds of validation through both human review and automatic MLLM verification. This rigorous process ensures the benchmark’s accuracy and reliability. By integrating these features, WorldSense sets a new standard for evaluating MLLMs’ ability to handle real-world complexity, bridging the gap between artificial intelligence and human-like multimodal understanding.
@@
We conduct comprehensive experiments on WorldSense to evaluate a range of existing MLLMs, including open-source video models, video-audio models, and proprietary models. Our results highlight significant limitations in the current models’ ability to understand and reason through omni-modal information in real-world contexts. First, while open-source video-audio models can process both video and audio inputs, they achieve only around 25% accuracy on our benchmark, comparable to random guessing. In contrast, proprietary models, such as Gemini 1.5 Pro, which processes both video and audio, achieve the highest accuracy at 48%. However, when provided with only a single modality (either video or audio), Gemini 1.5 Pro’s performance drops significantly by approximately 15%, underscoring the crucial role of omni-modal collaboration in real-world tasks. These findings emphasize the strong coupling of modalities in WorldSense and reveal a substantial gap in current MLLMs’ real-world understanding capabilities.
Additionally, we further conduct detailed ablation studies on how different modalities impact model performance. Our analysis shows that visual information plays significant for understanding, while adding audio—whether as raw signals or transcribed subtitles—also improves results. However, raw audio inputs yield superior gains, emphasizing the importance of paralinguistic features such as prosody, intonation, and acoustic context, which are often lost in text transcription. Additionally, increasing the density of video frame sampling, which provides richer temporal information, consistently improves performance. These findings underscore the complementary nature of visual and auditory modalities and highlight the necessity of integrating both for robust real-world understanding.
To summarise, in this paper, we have made three contributions: (i) we introduce, WorldSense, the first benchmark specifically designed to evaluate MLLMs’ omni-modal understanding capabilities in real-world scenarios. WorldSense is characterized by the integration of omni-modality, diverse video categories, and high-quality QA pairs; (ii) We have conducted extensive experiments to evaluate current MLLMs’ capabilities in real-world omni-modal understanding. Our results demonstrate that open-source video-audio MLLMs perform only marginally better than random guessing, while even the best-performing proprietary models achieve only accuracy, revealing significant room for accurate omni-modal understanding; (iii) Through detailed analysis, we investigate the crucial factors affecting real-world omni-modal understanding. Our findings demonstrate that acoustic information, visual cues, and dense temporal sampling all substantially impact model performance. These finds provide valuable guidance for the future development of real-world omni-modal understanding.
@@ 2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as [75, 35, 80, 11, 64, 51], successfully combine vision and text modalities. Subsequent research extends to temporal understanding [66, 24, 58, 39, 61, 65, 26, 14, 69, 74, 60, 6, 41, 34], while parallel efforts [55, 9, 8] focus on audio processing. Recently, researchers shift attention to models [7, 54, 58, 42, 59] capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
@@ 3 WorldSense
In this section, we aim to detail the construction of WorldSense, including data collection procedure, annotation pipeline, and statistics. Unlike existing benchmarks that assess modalities in isolation, WorldSense evaluates the MLLMs’ ability to perceive, understand, and reason about real-world scenarios through the integration of omni-modal information. As shown in Figure 1, all the multiple-choice questions are carefully crafted, to make sure the questions can only be answered through the comprehensive analysis of text, vision, and audio.
@@ 3.1 Design Principle
As for multi-modal evaluation, we base on the audio-visual synchronized videos, which capture temporal events, motion patterns, and audio-visual correlations. To curate the benchmark, we adhere to the following three principles, to ensure a rigorous and comprehensive evaluation.
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi-task Open-domain Sub./Aud. A-V Correlations MSRVTT-QA [68] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [72] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [30] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ Video-Bench [47] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [44] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [70] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [29] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [31] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [19] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [18] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
Comprehensive Domain. To ensure a thorough evaluation of MLLMs’ real-world understanding, we developed a systematic taxonomy that covers diverse domains and scenarios. This process started with primary categories reflecting core aspects of human experience, which were further subdivided into 67 subcategories to capture specific contexts. This hierarchical structure ensures that our video collection spans a broad range of real-world experiences, providing an ecologically valid basis for assessing multimodal understanding.
Diverse Acoustic Signals. In real-world scenarios, audio signals can be principally categorized into three fundamental types: speech, event and music. Our benchmark incorporates all three to ensure comprehensive coverage, enabling MLLMs to process and understand a wide spectrum of acoustic information—from semantic speech to abstract music and environmental sounds.
Multilevel Assessment. To evaluate the MLLMs’ perceptual and cognitive capabilities, we designed a multi-scale assessment at three levels: recognition (basic audiovisual element detection), understanding (comprehension of multimodal relationships), and reasoning (advanced cognitive tasks like causal inference and abstract thinking). We developed 26 tasks across to assess different aspects of multimodal comprehension, with a focus on integrating audio-visual information at each level.
@@ 3.2 Data Collection & Curation
We primarily source our video content from FineVideo [15], a large-scale dataset of high-quality YouTube videos with strong audio-visual correlations across diverse real-world scenarios. To enhance the coverage of musical content, we additionally incorporate videos from MusicAVQA [29].
Our data collection process involves a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, as illustrated in Figure 3(a). The process follows three key steps: (i) filtering videos based on predefined taxonomic categories described in Section 3.1 for comprehensive coverage; (ii) utilizing pre-computed metrics, including audio-visual correlation and dynamic content scores, to identify significant clips from an initial pool of approximately 8,000 videos; and (iii) conducting human expert review to assess video quality and real-world relevance. This rigorous process ultimately yields 1,662 high-quality video segments featuring strong audio-visual correlations across diverse real-world contexts.
@@ 3.3 Annotation Protocol
QA Annotation. A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs. For each video clip, annotators conduct thorough reviews of both visual and auditory content to ensure comprehensive understanding. They then generate questions and corresponding answer options that specifically require the integration of both visual and audio information for correct responses, thereby enabling effective evaluation of MLLMs’ multimodal understanding capabilities.
Quality Control. To ensure the quality of question-answer pairs, we implement a rigorous quality control process that combines human expertise with automated verification, as illustrated in Figure 3(b). Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) necessity of both visual and audio information for correct answers, and (iii) appropriateness of the question’s difficulty. Questions that fail to meet these standards are returned for revision.
We additionally employ MLLMs for automated verification. Vision-language models like Qwen2-VL[61] verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2[7] and OneLLM [21] are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
@@
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [42] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [42] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [42] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [21] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 Open-Source Video MLLMs Video-LLaVA [32] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [20] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [62] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [71] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [26] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [78] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [24] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [58] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
@@ 3.4 Dataset Statistics
As summarized in Table 1, the WorldSense dataset contains 1,662 video clips with synchronized audio, distributed across 8 primary categories and 67 subcategories. The average duration is 141.1 seconds, with lengths ranging from 30 seconds to over 10 minutes, capturing a wide variety of events and activities. In total, WorldSense includes 3,173 multiple-choice questions, covering three levels.
WorldSense encompasses diverse audio components including speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [19], OmniBench [31]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first benchmark designed to evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a more comprehensive benchmark for real-world multimodal understanding assessment.

### v2

@@ Abstract
We introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs. In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality. Based on our WorldSense, we extensively evaluate various state-of-the-art models. The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ( best accuracy). By analyzing the limitations of current models, we aim to provide valuable insight to guide development of real-world understanding. We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
@@ 1 Introduction
The ability to comprehend and reason about multimodal inputs—ranging from visual and textual to auditory, tactile, and beyond—is fundamental for both human and artificial agents to navigate and interpret the world. For example, when driving a car, a human driver integrates visual information (e.g., recognizing road signs, traffic lights, and obstacles), auditory cues (e.g., hearing the honking of another car or a siren approaching from behind), and tactile feedback (e.g., the feel of the steering wheel, the vibrations of the road, or the responsiveness of the brakes) to make real-time decisions and ensure safe navigation. This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and development of embodied agents designed to interact naturally.
@@
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [51, 25, 50, 59, 61, 79, 44, 12] have led to remarkable progress on a series of tasks, for example, classification [37], captioning [1, 10, 36], question-answering [58, 52, 41], OCR [47, 80], segmentation [26, 69, 23], autonomous driving [48, 54, 4] and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have started incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench [32] and AV-Odyssey Bench [20] mainly emphasize image evaluation, whereas other benchmarks [19, 30, 74] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
@@
This paper presents WorldSense, the first comprehensive benchmark designed to evaluate Multimodal Large Language Models (MLLMs) in perceiving, understanding, and reasoning with omni-modal information in real-world settings. The benchmark is defined by three key features: (i) Omni-modal integration. THe benchmark emphasizes the joint processing of audio and visual modalities, as illustrated in Figure 1. Each question requires both modalities for accurate response—removing either results in failure—enabling rigorous assessment of a model’s capacity for integrated sensory understanding. (ii) Diverse videos and task coverage. The benchmark includes 1,662 synchronized audio-visual videos spanning 8 domains and 67 fine-grained subcategories. It features 3,172 multiple-choice questions across 26 cognitive tasks, ranging from basic perception to high-level reasoning. This diversity supports systematic evaluation of multimodal comprehension across a broad task spectrum. (iii) High-quality annotations. All question-answer pairs are curated by 80 expert annotators and undergo multiple validation rounds, including human review and automated MLLM verification. This ensures annotation accuracy and benchmark reliability. Through these methodological advancements, WorldSense sets a new standard for evaluating MLLMs in real-world multimodal reasoning, advancing the field toward more human-like understanding.
@@
We conduct extensive evaluations for a broad spectrum of MLLMs, including open-source video models, video-audio models, and proprietary systems. Results reveal significant limitations in current models’ ability to reason over omni-modal inputs in real-world contexts. Specifically, open-source video-audio models, despite processing both modalities, achieve only 25% accuracy—comparable to random guessing. In contrast, proprietary models such as Gemini 1.5 Pro reach up to 48% accuracy. However, when restricted to a single modality (audio or video), Gemini’s performance drops by 15%, highlighting the critical role of integrated modality processing.
We further conduct ablation studies to dissect modality contributions. Visual inputs are essential, while audio—especially raw signals—yields additional gains over text transcriptions, due to preserved paralinguistic cues, e.g., prosody, intonation, acoustic context. These findings affirm the complementary nature of audio-visual information and the necessity of their joint modeling for robust real-world understanding. Failure case analysis reveals persistent limitations in current MLLMs, motivating future directions for improving multimodal reasoning.
To summarize, we have made the following contributions: (i) we present WorldSense, the first benchmark tailored for evaluating MLLMs’ ability on omni-modal video understanding, characterized by integrated audio-visual inputs, diverse content, and high-quality question-answering annotations; (ii) we have conducted extensive evaluation of existing MLLMs, showing that most open-source models perform near chance, and even the best proprietary model achieves only 48% accuracy—exposing a significant gap in real-world omni-modal reasoning; (iii) through ablation and failure analysis, we identify the key factors influencing performance, including raw audio and visual cues, and provide actionable insights to guide future omni-modal understanding design.
@@ 2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as [79, 36, 84, 11, 66, 53, 45, 12], successfully combine vision and text modalities. Subsequent research extends to temporal understanding [68, 25, 60, 40, 63, 67, 27, 14, 72, 78, 62, 6, 42, 35, 73], while parallel efforts [57, 9, 8] focus on audio processing. Recently, researchers shift attention to models [7, 56, 60, 43, 61] capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
@@ 3 WorldSense
In this section, we first introduce the design principles in Section 3.1, followed by a description of the data collection (Section 3.2) and annotation processes ( 3.3). We then compare statics of WorldSense with previous benchmarks in Section 3.4, and finally present our evaluation methodology 3.5.
@@ 3.1 Design Principle
As for multi-modal evaluation, we base on the audio-visual synchronized videos, which capture temporal events, motion patterns, and audio-visual correlations. To curate the benchmark, we adhere to the following three principles, to ensure rigorous and comprehensive evaluations for MLLMs.
Comprehensive Domain Coverage. To capture the diversity of real-world scenarios, we construct a hierarchical taxonomy starting from broad human-centric domains, refined into 67 fine-grained subcategories. This structure ensures wide ecological coverage, enabling robust assessment of multi-modal understanding across varied contexts.
Diverse Acoustic Modalities. Real-world audio can be broadly classified into speech, environmental events, and music. The benchmark includes all three types, enabling evaluation across a spectrum of acoustic complexity—from linguistic content to non-verbal and abstract auditory cues.
Multilevel Cognitive Assessment. We design a three-tiered evaluation framework targeting: recognition (detection of basic audio-visual elements), understanding (comprehension of multimodal relationships), and reasoning (high-level inference tasks such as causal inference or abstract thinking). The benchmark includes 26 tasks aligned with these levels, encouraging holistic evaluation of perceptual and cognitive capabilities in multimodal settings.
@@ 3.2 Data Collection & Curation
We primarily source our video content from FineVideo [15], a large-scale dataset comprising high-quality YouTube videos that exhibit strong audio-visual correlations across diverse real-world scenarios. To enrich the benchmark’s coverage of musical content, we supplement it with selected videos from MusicAVQA [30], ensuring a more balanced representation of auditory modalities.
Our data collection employs a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, following three main steps in Figure 3(a): (i) filtering videos according to predefined taxonomic categories delineated in Section 3.1; (ii) selecting clips based on pre-computed audio-visual correlation and dynamic content metrics from about 8,000 initial videos; and (iii) human expert review for video quality and real-world relevance. This rigorous selection and processing results in 1,662 high-quality video segments with strong audio-visual correlations across various real-world scenarios.
@@ 3.3 Annotation Protocol
Question-Answering (QA) Annotation. A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs for each video by thoroughly reviewing both visual and audio content. The questions are designed to require integration of multiple modalities, enabling effective assessment of MLLMs’ multimodal understanding.
Quality Control. To ensure QA quality, we implement a rigorous quality control process combining expert review and automated checks, as illustrated in Figure 3(b). Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) multimodal necessity for correct answers, and (iii) appropriate difficulty. Questions that fail to meet these standards are returned for revision.
We also use MLLMs for automated verification. Vision-language models like Qwen2-VL[63] verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2[7] and OneLLM [22] are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
@@ 3.4 Dataset Statistics
As summarized in Table 1, our proposed WorldSense benchmark contains 1,662 video clips with synchronized audio across 8 categories and 67 subcategories, averaging 141.1 seconds in length, including 3,173 multiple-choice questions on three cognitive levels.
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi task Open domain Sub./ Aud. A-V Correlations MSRVTT-QA [70] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [76] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [31] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ MovieChat [55] V 130 500.0 1,950 M - ✗ ✗ ✓ ✗ ✗ Video-Bench [49] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [46] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [74] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [30] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [32] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [20] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [19] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
WorldSense features diverse audio types such as speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [20], OmniBench [32]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first to comprehensively evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a comprehensive benchmark for real-world multimodal understanding assessment.